# Bipartite Graph Variational Auto-Encoder with Fair Latent Representation to Account for Sampling Bias in Ecological Networks

### v2

@@ 3 Bipartite and fair graph variational auto-encoder
where $\delta$ is a hyperparameter, $KL$ is the Kullback-Leibler divergence, and $p_{1},p_{2}$ are Gaussian priors for $Z_{1}$ and $Z_{2}$. This method can also be extended to the case where we also seek independence between $Z_{2}$ and another protected variable.
3.1 Model adaptation to the Spipoll data set
3.1.1 Paradigm shift
@@
We applied our methodology on the Spipoll data set. We extracted covariates on climatic conditions at each date and place of observations using the European Copernicus Climate data set provided by Cornes et al. (2018). A common practice in ecology to study plant-pollinator interactions is to consider plant and insect species as nodes of a bipartite network, with edges determined by the interactions observed between the two. In the case of Spipoll data set, this implies aggregating all observations of the interactions between a given plant and insect species but also all the covariates describing observation conditions, such as date and climate. Such aggregation is not straightforward so we propose to change paradigm by considering a bipartite network where the first type of nodes is the session of observations, and the second type corresponds to insects observed during the session. Each session has the previously mentioned covariates and a one-hot encoding describing the plant genus. This paradigm allows directly using the Spipoll data without doing any aggregation. Link prediction task in this situation aims to predict which insect will be present during a given observation session. However, we still wish to ultimately obtain a bipartite plant-insect network, as this is the most widely used tool in this field of study. To ensure that the latent space could also be used to create a plant-insect network, we propose two methods to generate the corresponding plant-insect network, using the same decoder. We propose in our setting to define $S$, the protected variable, as the number of participation from the user. This number of participation would work as a proxy of the user’s experience. By employing this measure, we aim to construct a latent space that remains unaffected by variations in observers’ experience levels.
3.1.2 Model
@@
In this part, we elaborate on the application of our methodology, taking into account the specific requirements of the Spipoll data set. We consider $B$ our $n_{1}\times n_{2}$ incidence matrix, where the $n_{1}$ rows correspond to the number of sampling session, and the $n_{2}$ columns correspond to the number of different observed pollinators in the dataset. For all $i$ and $j$ , $B_{i,j}\in\{0,1\}$ describes the absence or the presence of the pollinator $j$ during the session $i$. Let $P=(P_{i,k})$, $i=1,\dots,n_{1},k=1,\dots,u$ where $u$ corresponds to the number of observed taxa of plants. $P_{i,k}\in\{0,1\}$ is a binarized categorical variable that describes the plant taxonomy of the $i^{th}$ session. For all $i$, there is only one coordinates $k$ such that $P_{i,k}=1$ while the others are equal to 0. To build the binary adjacency matrix $B^{\prime}$ of plant-pollinator interactions from the session-pollinator matrix $B$, we compute $B^{\prime}=\mathbf{1}(P^{\top}B>0)$. We create a second latent space from a realization of the first:
@@
where $Z^{\prime}_{1}$ is a $u\times D$ matrix, the rows of which $(Z^{\prime}_{1l}\in\mathbb{R}^{D})_{1\leq l\leq u}$ are the latent variables associated to each plant. For any $l$, $\tilde{q}_{1}(Z^{\prime}_{1l}|Z_{1},P)$ samples uniformly one $Z_{1i}$ among the ones where $P_{i,l}=1$. We use the same decoder for both latent space. Let $\tilde{P}_{i,k}=\frac{P_{i,k}}{\sum_{l=1}^{n_{1}}P_{l,k}}.$ It is also possible to estimate $B^{\prime}$ from the reconstruction $\widehat{B}$ itself, by calculating $\widehat{B^{\prime}}=\tilde{P}^{\top}\widehat{B}$. This is equivalent to calculate the mean by plants of the predicted probabilities of interaction. The model is summarized in Figure 2.
[…]
@@
The protected variable $S$ is the log base 10 of the number of observation session the user has already performed. To fit this model, we minimize the same loss as in Section 3 with a supplementary term
@@ 4 Application on simulated data
This simulation study tries to replicate numerically the sampling process taking place in the Spipoll data set, and to see if our method can yield better result than not taking into account the user’s experience. We also compare our method with a standard adversarial learning. An application of the fair and bipartite graph variational auto-encoder on a simple simulated case and the definition of the adversarial methodology (Zhang et al., 2018) are available in the supplementary materials, Section B.1
@@ 4.1 Settings
Session-pollinator network : Let $n_{1}=3000$ be the number of observers, we suppose for the simulation that their user’s experience is given for $1\leq i\leq n_{1}$ by $S_{i}=round(S^{\prime}_{i})+1,$ where $S^{\prime}_{i}\sim\varepsilon(21)$, with $\varepsilon(\lambda)$ an exponential random variable of density $f_{\lambda}(x)=\frac{1}{\lambda}e^{-\frac{x}{\lambda}}\mathbf{1}_{\{x\geq 0\}}$. Each user will select uniformly at random one flower $k$, and will observe $obs_{i}=round(2\log(S_{i}))$ possible interactions from the $k$-th row of the matrix $B_{0}^{\prime}$ at random, with a probability proportional to the difficulty, given by the weight $w_{easy}$ and $w_{hard}$. Once the observations-insects network is constructed, we also have access to the observed plant-pollinator network. A summary of the procedure is available in Figure 3.
@@ 4.2 Results
Table 1: Comparison between the Bipartite variational graph auto-encoder and its fair counterparts with simulated data. In the table are reported the mean and standard deviation for 30 trials, except for #$\#p_{0.05}$ which is a count.
@@
AUC($\widehat{B}$) $0.641\pm 0.013$ $0.595\pm 0.021$ $0.642\pm 0.014$
AUC($\tilde{B}^{\prime}$) $0.582\pm 0.017$ $0.592\pm 0.024$ $0.580\pm 0.021$
AUC($\widehat{B^{\prime}}$) $0.613\pm 0.020$ $0.655\pm 0.040$ $0.608\pm 0.026$
@@
The results for the link prediction task for the simulated data set are summarized in Table 1. The predictions were made with the complete dataset $B$, that has been split with 30% of the edges hidden. 20% of these hidden edges are used as validation data set, and the remaining 10% for the test set. Both sets also contain an equivalent amount of non-edges that are not in the train set. Moreover, as we also try to have a latent representation for $B^{\prime}$, we carefully construct another training set with the edges and non-edges that have already been used for the learning of $B$. The remaining edges and non-edges are used for the test, from which we know the ground truth thanks to the true underlying matrix $B^{\prime}_{0}$. We compare the link prediction $\widehat{B}$ at the observation session level (AUC($\widehat{B}$)), and the link prediction $\widehat{B^{\prime}}$ and $\tilde{B^{\prime}}$ at the plant-insect level (AUC($\widehat{B^{\prime}}$), AUC($\tilde{B^{\prime}}$)). AUC($\tilde{B^{\prime}}$) is calculated with the prediction obtained by the second latent space $Z^{\prime}_{1},Z_{2}$, whereas AUC($\widehat{B^{\prime}}$) is calculated from the prediction obtained by the reconstruction $\widehat{B}$ with $\widehat{B^{\prime}}=\tilde{P}^{\top}\widehat{B}$ (see Figure 2). We also calculate the Spearman correlation $cor_{sp}(\widehat{B^{\prime}})$ between the predicted matrix $\widehat{B^{\prime}}$ and the true underlying matrix of probabilities given by the SBM. Both methods are fit with 1 000 iterations of Adam algorithm with learning rate 0.01, and with a latent space of dimension 4 with $D_{+}=D_{-}=2$, using a computer equipped with an Intel Xeon(R) CPU E5-1650 v4 and 32GB of RAM.
@@
From Table 1, we notice a reduction of prediction accuracy for $B$ when enforcing a fair setting in the BVGAE. This is a common consequence of most fairness settings. However, using fairness in this case has increased the quality of prediction at the plant-pollinator level ($B^{\prime}$) as we can see with the AUC($\widehat{B^{\prime}}$) and the Spearman correlation obtained with the fair-BVGAE. The reconstruction of the plant-pollinator level is clearly better via $\widehat{B^{\prime}}$ than via $\tilde{B^{\prime}}$. Moreover, the latent space given by the BVGAE is not independent of the protected variable $S$. This can be seen by looking at the p-value of the HSIC independence test and the correlation between $\tilde{Z_{1}}$ and $S$. Even if it is not enough to guarantee independence, we can see that the correlation between the latent space and the protected variable is much higher in the BVGAE than in the fair-BVGAE and ADV model. In all the simulations, the independence hypothesis has been rejected for the BVGAE and kept for the fair-BVGAE. The ADV did not manage to have a smaller HSIC than the BVGAE, and the independence hypothesis was rejected most of the time. The ADV model is much harder to calibrate because it requires a second neural network to optimize.
@@ 5 Results on the Spipoll data set
AUC($\widehat{B}$) $0.869\pm 0.003$ $0.834\pm 0.007$
AUC($\tilde{B^{\prime}}$) $0.731\pm 0.004$ $0.710\pm 0.016$
AUC($\widehat{B^{\prime}}$) $0.777\pm 0.018$ $0.758\pm 0.022$
@@
We consider the observation period of the Spipoll data set from 2017 to 2020 included, in metropolitan France and Belgium. We consider a total of $n_{1}=12754$ observation sessions, where $n_{2}=306$ taxa of insects and $u=83$ genus of plants have been observed. The observation session-insect matrix $B$ has a total of 94 909 interactions reported, and the plant-insect matrix $B^{\prime}$ has 9 754 different interactions. The results for the link prediction task for the Spipoll data set are summarized in Table 2. We use the same training setting as previously, with $D_{+}=D_{-}=2$. Contrary to the simulation study, the underlying plant-pollinator interaction matrix $B^{\prime}_{0}$ is not available. Therefore, the AUC for the prediction of $B^{\prime}$ are performed on the observed plant-pollinator network which is incomplete.
@@
Looking at the results in Table 2, both methods have better prediction on $B$ than on $B^{\prime}$, and calculating $\widehat{B^{\prime}}=\tilde{P}^{\top}\widehat{B}$ yields better results than calculating $\tilde{B^{\prime}}$. The BVGAE has better AUC than its fair-counterpart, however, even if the linear correlation between the latent space and the protected variable $S$ seems low (0.051), the latent space of the BVGAE is not independent of $S$ according to the HSIC test. Although the AUC($\widehat{B}$) decreased in average from 0.869 to 0.834, the fair-BVGAE has a latent space independent of $S$, which is the target result: the latent representation $Z_{1}$ is independent of the users’ experience levels. This can yield better prediction as seen in our simulation study, but it is difficult to assert because the true underlying network is inaccessible. Looking at Figure 4 we can see a change of structure when accounting for sampling bias leading to a change of our ecological understanding of this plant-pollinator network. For example, the fair adjustment overall increases the probability of connection, notably of Lepidoptera revealing a higher contribution of butterflies to pollination. A checkerboard structure is observed for the interactions of Hymenoptera in the BVGAE suggesting contrasted preferences among plants and pollinators. This structure is smoothened in the fair-BVGAE indicating that such differences in preferences have no ecological ground but are related to the sampling process. In addition to the plant-pollinator network, the method provides a session-pollinator embedding, the ecological interest of which is demonstrated in Section C.2. Full representations of latent spaces are available in Section C.1.

### v3

@@ 2.4 Bipartite and fair graph variational auto-encoder
where $\delta$ is a hyperparameter, $KL$ is the Kullback-Leibler divergence, and $p_{1},p_{2}$ are Gaussian priors for $Z_{1}$ and $Z_{2}$. This method can also be extended to the case where we also seek independence between $Z_{2}$ and another protected variable. To have a better understanding of the methodology, we can draw a parallel in A with a linear embedding for data that are elements of a real vector space. We can reformulate the problem of principal component analysis under independence constraint as an optimization problem with a supplementary term that encourages the independence.
3 Model adaptation to the Spipoll data set
3.1 Paradigm shift
A common practice in ecology to study plant-pollinator interactions is to consider plant and insect species as nodes of a bipartite network, with edges determined by the interactions observed between the two. In the case of Spipoll data set, this implies aggregating all observations of the interactions between a given plant and insect species but also all the covariates describing observation conditions, such as date and climate.
@@ 3.1 Paradigm shift
Such aggregation is not straightforward so we propose to change paradigm by considering a bipartite network where the first type of nodes is the session of observations, and the second type corresponds to insects observed during the session. Each session has the previously mentioned covariates and a one-hot encoding describing the plant genus. This paradigm allows directly using the Spipoll data without doing any aggregation. Link prediction task in this situation aims to predict which insect will be present during a given observation session. However, we still wish to ultimately obtain a bipartite plant-insect network, as this is the most widely used tool in this field of study. To ensure that the latent space could also be used to create a plant-insect network, we propose two methods to generate the corresponding plant-insect network, using the same decoder. We propose in our setting to define $S$, the protected variable, as the number of participation from the user. This number of participation would work as a proxy of the user’s experience. By employing this measure, we aim to construct a latent space that remains unaffected by variations in observers’ experience levels.
3.2 Model
@@
In this part, we elaborate on the application of our methodology, taking into account the specific requirements of the Spipoll data set. We consider $B$ our $n_{1}\times n_{2}$ incidence matrix, where the $n_{1}$ rows correspond to the number of sampling session, and the $n_{2}$ columns correspond to the number of different observed pollinators in the dataset. For all $i$ and $j$ , $B_{i,j}\in\{0,1\}$ describes the absence or the presence of the pollinator $j$ during the session $i$. Let $P=(P_{i,k})$, $i=1,\dots,n_{1},\ k=1,\dots,u$ where $u$ corresponds to the number of observed taxa of plants. $P_{i,k}\in\{0,1\}$ is a binarized categorical variable that describes the plant taxonomy of the $i^{th}$ session. For all $i$, there is only one coordinate $k$ such that $P_{i,k}=1$ while the others are equal to 0. To build the $u\times n_{2}$ binary adjacency matrix $B^{\prime}$ of plant-pollinator interactions from the session-pollinator matrix $B$,we check for each plant $k$ and each pollinator $j$ if there are any session $i$ where $P_{i,k}=1$ and $B_{i,j}=1$. Should it be the case, then $B^{\prime}_{k,j}=1$, otherwise $B^{\prime}_{k,j}=0$. This can be computed using $B_{k,j}^{\prime}=\mathbf{1}\{(P^{\top}B)_{k,j}>0\}$, where $\mathbf{1}$ is the indicator function.
We create a second latent space from a realization of the first:
@@
where $Z^{\prime}_{1}$ is a $u\times D$ matrix, the rows of which $(Z^{\prime}_{1l}\in\mathbb{R}^{D})_{1\leq l\leq u}$ are the latent variables associated to each plant. For any $l$, $\tilde{q}_{1}(Z^{\prime}_{1l}|Z_{1},P)$ samples uniformly one $Z_{1i}$ among the ones where $P_{i,l}=1$. We use the same decoder for both latent space.
It is also possible to estimate $B^{\prime}$ from the reconstruction $\widehat{B}$ itself, by averaging all the predicted probabilities of interaction by plants. This average is estimated using $\tilde{P}_{i,k}=\frac{P_{i,k}}{\sum_{l=1}^{n_{1}}P_{l,k}}$ and $\widehat{B^{\prime}}=\tilde{P}^{\top}\widehat{B}$.
$B,X_{1},X_{2}$$Z_{1},Z_{2}$$^{\widehat{B}}$$\text{AUC}(\widehat{B},B)$$B^{\prime}$$^{\widehat{B^{\prime}}}$$\text{AUC}(\widehat{B^{\prime}},B^{\prime})$$Z^{\prime}_{1},Z_{2}$$\tilde{B^{\prime}}$$\text{AUC}(\tilde{B^{\prime}},B^{\prime})$$\scriptstyle{q}$$\scriptstyle{\mathbf{1}\{(P^{\top}B)>0\}}$$\scriptstyle{p}$$\scriptstyle{\tilde{q}}$$\scriptstyle{\tilde{P}^{T}\widehat{B}}$$\scriptstyle{p}$
@@
The protected variable $S$ is the log base 10 of the number of observation session the user has already performed. To fit this model, we minimize the same loss as in Section 2.4 with a supplementary term
@@ 4 Application on simulated data
This simulation study tries to replicate numerically the sampling process taking place in the Spipoll data set, and to see if our method can yield better result than not taking into account the user’s experience. ,We also compare our methodology with an adversarial learning algorithm (ADV) (zhangMitigatingUnwantedBiases2018) where the output $\mu_{1}$ is then used as an input to a 4-layer perceptron, which attempts to predict the protected variable $S$. The loss is then penalized if the predicted output is correlated with the protected variable. An application of the fair and bipartite graph variational auto-encoder on a simple simulated case is available in the supplementary materials, B.1.
@@ 4.1 Settings
Session-pollinator network : Let $n_{1}=3000$ be the number of observers, we suppose for the simulation that their user’s experience is given for $1\leq i\leq n_{1}$ by $S_{i}=round(S^{\prime}_{i})+1,$ where $S^{\prime}_{i}\sim\varepsilon(21)$, with $\varepsilon(\lambda)$ an exponential random variable of density $f_{\lambda}(x)=\frac{1}{\lambda}e^{-\frac{x}{\lambda}}\mathbf{1}_{\{x\geq 0\}}$. Each user will select uniformly at random one flower $k$, and will observe $obs_{i}=round(2\log(S_{i}))$ possible interactions from the $k$-th row of the matrix $B_{0}^{\prime}$ at random, with a probability proportional to the difficulty, given by the weight $w_{easy}$ and $w_{hard}$. Once the observations-insects network is constructed, we also have access to the observed plant-pollinator network. The covariates related to the observation sessions are defined by $X_{1}=P$ where $P$ is the binarized categorical variable (83 columns) describing the observed plant genus. No feature is considered for the pollinators, therefore $X_{2}$ is set as the identity matrix of $n_{2}$ rows, as done in Kipf. A summary of the procedure is available in Figure 3.
@@ 4.2 Results
Table 1: Comparison between the Bipartite variational graph auto-encoder and its fair counterparts with simulated data. In the table are reported the mean and standard deviation for 30 trials, except for $p_{0.05}$ which is a count.
@@
AUC($\widehat{B},B$) $0.641\pm 0.013$ $0.595\pm 0.021$ $0.642\pm 0.014$
AUC($\tilde{B}^{\prime},B^{\prime}_{0}$) $0.582\pm 0.017$ $0.592\pm 0.024$ $0.580\pm 0.021$
AUC($\widehat{B^{\prime}},B^{\prime}_{0}$) $0.613\pm 0.020$ $0.655\pm 0.040$ $0.608\pm 0.026$
@@
The results for the link prediction task for the simulated data set are summarized in Table 1. The predictions were made with the complete dataset $B$, that has been split with 30% of the edges hidden. 20% of these hidden edges are used as validation data set, and the remaining 10% for the test set. Both sets also contain an equivalent amount of non-edges that are not in the train set. Moreover, as we also try to have a latent representation for $B^{\prime}$, we carefully construct another training set with the edges and non-edges that have already been used for the learning of $B$. The remaining edges and non-edges are used for the test, from which we know the ground truth thanks to the true underlying matrix $B^{\prime}_{0}$. We compare the link prediction $\widehat{B}$ at the observation session level (AUC($\widehat{B},B$)), and the link prediction $\widehat{B^{\prime}}$ and $\tilde{B^{\prime}}$ at the plant-insect level (AUC($\widehat{B^{\prime}},B^{\prime}_{0}$), AUC($\tilde{B^{\prime}},B^{\prime}_{0}$)). AUC($\tilde{B^{\prime}},B^{\prime}_{0}$) is calculated with the prediction obtained by the second latent space $Z^{\prime}_{1},Z_{2}$, whereas AUC($\widehat{B^{\prime}},B^{\prime}_{0}$) is calculated from the prediction obtained by the reconstruction $\widehat{B}$ with $\widehat{B^{\prime}}=\tilde{P}^{\top}\widehat{B}$ (see Figure 2). We also calculate the Spearman correlation $cor_{sp}(\widehat{B^{\prime}})$ between the predicted matrix $\widehat{B^{\prime}}$ and the true underlying matrix of probabilities given by the SBM. Both methods are fit with 1 000 iterations of Adam algorithm with learning rate 0.01, and with a latent space of dimension 4 with $D_{+}=D_{-}=2$, using a computer equipped with an Intel Xeon(R) CPU E5-1650 v4 and 32GB of RAM.
@@
From Table 1, we notice a reduction of prediction accuracy for $B$ when enforcing a fair setting in the BVGAE. This is a common consequence of most fairness settings. However, using fairness in this case has increased the quality of prediction at the plant-pollinator level ($B^{\prime}$) as we can see with the AUC($\widehat{B^{\prime}},B^{\prime}_{0}$) and the Spearman correlation obtained with the fair-BVGAE. The reconstruction of the plant-pollinator level is clearly better via $\widehat{B^{\prime}}$ than via $\tilde{B^{\prime}}$.
Moreover, the latent space given by the BVGAE is not independent of the protected variable $S$. This can be seen by looking at the p-value of the HSIC independence test and the correlation between $\tilde{Z_{1}}$ and $S$. Even if it is not enough to guarantee independence, we can see that the correlation between the latent space and the protected variable is much higher in the BVGAE than in the fair-BVGAE and ADV model. In all the simulations, the independence hypothesis has been rejected for the BVGAE and kept for the fair-BVGAE. The ADV did not manage to have a smaller HSIC than the BVGAE, and the independence hypothesis was rejected most of the time. The ADV model is much harder to calibrate because it requires a second neural network to optimize.
@@ 5 Results on the Spipoll data set
AUC($\widehat{B},B$) $0.869\pm 0.003$ $0.834\pm 0.007$
AUC($\tilde{B^{\prime}},B^{\prime}$) $0.731\pm 0.004$ $0.710\pm 0.016$
AUC($\widehat{B^{\prime}},B^{\prime}$) $0.777\pm 0.018$ $0.758\pm 0.022$
@@
We consider the observation period of the Spipoll data set from 2017 to 2020 included, in metropolitan France and Belgium. We consider a total of $n_{1}=12754$ observation sessions, where $n_{2}=306$ taxa of insects and $u=83$ genera of plants have been observed. The covariates related to the observation sessions are $X_{1}=[P,T]$ where $P$ is the binarized categorical variable (83 columns) describing the observed plant genus, concatenated with $T$, the average temperature on the day at the observation location, provided by the European Copernicus Climate data (cornesEnsembleVersionEOBS2018). No feature is considered for the pollinators, therefore $X_{2}$ is set as the identity matrix with $n_{2}$ rows. We use the same training setting as previously, with $D_{+}=D_{-}=2$. We justify this choice by looking at the estimated mean of AUC(B^)AUC($\widehat${B}) for different numbers of dimensions $D_{+}$ and $D_{-}$ in C.1. Contrary to the simulation study, the underlying plant-pollinator interaction matrix $B^{\prime}_{0}$ is not available. Therefore, the AUC for the prediction of $B^{\prime}$ are performed on the observed plant-pollinator network which is incomplete.
@@
Looking at the results in Table 2, both methods have better prediction on $B$ than on $B^{\prime}$, and calculating $\widehat{B^{\prime}}=\tilde{P}^{\top}\widehat{B}$ yields better results than calculating $\tilde{B^{\prime}}$. The BVGAE has better AUC than its fair-counterpart, however, even if the linear correlation between the latent space and the protected variable $S$ seems low (0.051), the latent space of the BVGAE is not independent of $S$ according to the HSIC test. Although the AUC($\widehat{B}$) decreased in average from 0.869 to 0.834, the fair-BVGAE has a latent space independent of $S$, which is the target result: the latent representation $Z_{1}$ is independent of the users’ experience levels. This can yield better prediction as seen in our simulation study, but it is difficult to assert because the true underlying network is inaccessible. Looking at Figure 4 we can see a change of structure when accounting for sampling bias leading to a change of our ecological understanding of this plant-pollinator network. For example, the fair adjustment overall increases the probability of connection, notably of Lepidoptera revealing a higher contribution of butterflies to pollination. A checkerboard structure is observed for the interactions of Hymenoptera in the BVGAE suggesting contrasted preferences among plants and pollinators. This structure is smoothened in the fair-BVGAE indicating that such differences in preferences have no ecological ground but are related to the sampling process.
Figure 5: Focus on the embeddings provided by the fair-BVGAE for the observation sessions performed on the genera Daucus, Leucanthenmum and Lavandula.
In addition to the plant-pollinator network, the method provides a session-pollinator embedding. In order to analyze the pattern of plant pollinator interaction, we can focus on the sessions where a particular genera has been observed, in order to have a closer look on specific interactions or to compare interaction profiles. As an example in Figure 5, we focus on three plant genera, Daucus, Leucanthemum and Lavandula. We remind that the method provides an embedding of dimension $D=D_{+}+D_{-}=4$ with $D_{+}=D_{-}=2$, which means that for the first two dimensions, insects and sessions that are embedded in the same direction are more likely to be connected, and the ones in the opposite direction are less likely to be connected. On the contrary, insects and sessions that are embedded in the same direction for the third and fourth dimensions are less likely to be connected, while the ones in the opposite direction are more likely to be connected. In the top-left panel of Figure 5, sessions conducted on Daucus are positioned further to the right along the first axis compared to those conducted on Leucanthemum or Lavandula. Since most observed insects are also located on the right side of the plot, this suggests a higher probability of insect observations during sessions on Daucus than on the other genera and thereby that it is a more generalist plant than the two others. This can also be seen in the bottom-right panel, where sessions conducted on Lavandula are positioned in the same direction as most insects along the third and fourth axis, which suggests that Lavandula is more specialized than the other two genera in terms of plant–pollinator interactions. Representations of latent spaces with all plant genera are available in C.